OpenStudy (anonymous):

Solve the given problem. The ages of three sisters are three even consecutive integers. If the sum of the 1st, four times the 2nd, and twice the 3rd is 86, what are the three ages?

OpenStudy (anonymous):

x + 4(x+2) + 2(x+4) = 86 Solve for x.

OpenStudy (anonymous):

x+4x+8+2x+8=86 7x+16=86 7x=70 thus, x=10

OpenStudy (anonymous):

thus, the ages of 3 sis were 10 , 12 , & 14